OpenStudy (anonymous):

find the value of x if line RS bisects AB and RS = 36

OpenStudy (anonymous):

 This is the picture of what comes with the question. PLEASE HELP ME

OpenStudy (agent0smith):

Okay now i can see it better.. nm. If RS = 36, then 18+2y+6=36 find y first from that. Then we'll find x.

OpenStudy (anonymous):

y=12

OpenStudy (agent0smith):

2y=12

OpenStudy (anonymous):

y=6

OpenStudy (agent0smith):

Yep. Now since AB is bisected it's cut in half... so each of the halves are equal 25-3x=2y+5 and you know y=6 so you can find x.

OpenStudy (anonymous):

okay final answer is -8??? im sorry im probably annoying the crap out of you

OpenStudy (agent0smith):

lol no you're not. plug in y=6: 25-3x=2*6+5 now try solving this for x

OpenStudy (anonymous):

x=8 over 3 or 2.6 repeating??

OpenStudy (agent0smith):

Yep, looks correct.

OpenStudy (anonymous):

THANK YOU SO MUCH! have a great nightt